1/cos1*cos2+1/cos2*cos3+1/cos3*cos4+..............+1/cos88*cos89=?

Grade:12

1 Answers

sunil raikwar
10 years ago
1/sin1[sin(2-1)/cos1*cos2+sin(3-2)/cos2*cos3+..........sin(89-88)/cos88cos89] =1/sin1[tan2-tan1+tan3-tan2+tan4-tan3+.........tan89-tan88] =1/sin1[tan89-tan1] =1/sin1[sin89/cos89 -sin1/cos1] =1/sin1[sin(89-1)/cos1cos89] =1/sin1[sin88/cos1cos89] =2sin88/2sin1cos1.sin1 =2sin88/sin2.sin1 =2sin88/cos88.sin1 =2tan88.cosec1